# Higher order linear differential equations for unitary matrix integrals: applications and generalisations

### v1

@@ 4.1 The case of general parameters
Next we would like to relate (3.10) to a scalar linear differential equation of degree $n+1$ for $\hat{I}_{n}(x)$ (and thus by the replacement $b\mapsto b-1$ for $\hat{I}_{0}(x)$; recall (3.12)).
@@
Let $\hat{I}_{n}(x)$ be specified via the differential-difference system (3.8). For some polynomials $\{q_{m}(x)\}_{m=0,1,\dots,n-1}$ we have
$\Big{(}x^{n}{d^{n+1}\over dx^{n+1}}+q_{n}(x)x^{n-1}{d^{n}\over dx^{n}}+q_{n-1}(x)x^{n-2}{d^{n-1}\over dx^{n-1}}+\cdots+q_{2}(x)x{d^{2}\over dx^{2}}+q_{1}(x){d\over dx}+q_{0}(x)\Big{)}\hat{I}_{n}(x)=0.$ (4.1)
The degree in $x$ of the coefficient of each $d^{j}\over dx^{j}$} is [(n1+j)/2][(n-1+j)/2], thus restricting the degrees of the $q_{m}(x)$. Furthermore, when $\{\tilde{B}_{j}\}$ are all integers, the coefficients of these polynomials are also integers.
@@
Let $D_{x}:=x{d\over dx}$. From (3.8) in the case $p=n$ we read off that
$n\hat{I}_{n-1}(x)=(\tilde{B}_{n}+D_{x})\hat{I}_{n}(x).$ (4.2)
@@
$n(n-1)(n-2)\hat{I}_{n-3}(x)=\bigg{(}(\tilde{B}_{n-2}+D_{x})\Big{(}(\tilde{B}_{n-1}+D_{x})(\tilde{B}_{n}+D_{x})-nx\Big{)}-2x(n-1)(\tilde{B}_{n}+D_{x})\bigg{)}\hat{I}_{n}(x).$ (4.4)
@@
$x{d\over dx}+c_{0},\>\>x^{2}{d^{2}\over dx^{2}}+d_{0}x{d\over dx}+(c_{1}+c_{2}x),\>\>x^{3}{d^{3}\over dx^{3}}+e_{0}x^{2}{d^{2}\over dx^{2}}+(d_{1}+d_{2}x)x{d\over dx}+(c_{3}+c_{4}x).$ (4.5)
Repeating this procedure down to the case $p=1$ of (3.8) allows each $\hat{I}_{j}(x)$ ($j=0,\dots,n)$ to be written as a differential operator of degree njn-j acting on $\hat{I}_{n}(x)$. In particular, $n!\hat{I}_{1}(x)$ has the structure of a differential operator of the form (4.1) with $n\mapsto n-1$ and an extra factor of $x$ multiplying each of the terms involving a derivative. Similarly, $n!\hat{I}_{0}(x)$ has the structure of a differential operator of the form in (4.1) as written, except for an extra factor of $x$ multiplying each of the terms involving a derivative.
@@
${d\over dx}\hat{I}_{0}(x)-n\hat{I}_{1}(x)=0.$
From the established formulas for $\hat{I}_{0}(x),\hat{I}_{1}(x)$, this implies (4.1). That the coefficients of the polynomials $\{q_{j}(x)\}$ are integers for $\{\tilde{B}_{j}\}$ integers is immediate from the structure of the RHS of (4.2)–(4.4) as extends to $\hat{I}_{0}(x),\hat{I}_{1}(x)$. ∎
Remark 4.1.
The restriction on the degrees of the polynomials $\{q_{j}(x)\}$ relates to a conjecture from [10]. See too the discussion in [9].
@@
In the case $\beta=2$, $n=l$ and $\tilde{B}_{p}=p(l-p)$, $\hat{I}_{n}(x)$ is according to (2.9) and (3.13) (the latter modified in keeping with Remark 3.1.1) equal to the matrix average in (1.3) with $z=x^{1/2}$ and thus is a generating function for the sequence $\{T_{l}(N)\}_{N=1,2,\dots}$. Using different methods, the precise form of the corresponding differential operator in (4.1) has been computed for $l=2,3,\dots,7$ in [9]. From the substitution method given in the proof of Proposition 3.1 we can reclaim the listing of [9], and extend it to higher values of $l$ (see the work [41] for other applications in random matrix theory of this method in reducing first order linear matrix differential equations to higher order scalar differential equations). Here we make note of the case $l=8$.
@@
$q_{8}(x)=120,\>\>\>q_{7}(x)=3(1477-40x),\>\>\>q_{6}(x)=2(48809-4467x),$
$q_{5}(x)=6(203139-42433x+728x^{2}),\>\>\>q_{4}(x)=4(2177225-881163x+51840x^{2}),$
$q_{3}(x)=4(8411053-6187600x+856062x^{2}-13120x^{3}),$
$q_{2}(x)=(7322954-10406303x+2993572x^{2}-161622x^{3}),$
$q_{1}(x)=16(1456832-6870928x+4305495x^{2}-584784x^{3}+9216x^{4}),$
$q_{0}(x)=32(-728416+1821603x-643914x^{2}+36864x^{3}).$
@@
In relation to this characterisation for general $l$, substituting the series on the RHS of (1.3) transforms the scalar linear differential equation to a scalar linear difference equation with coefficients that are polynomials in $N$, and of degree no greater than $l+1$; see [9] for more on this and [11, footnote 22] for references containing the explicit forms up to $l=5$. However, in relation to computing the power series solution, there is no advantage in first reducing the matrix differential equation down to a scalar differential equation, rather use of the analogue of (3.18) (which is based on the matrix differential equation) is more efficient.
On this last point, it is of interest to compare our scheme for the computation of $\{T_{l}(N)\}$, which we will base on (3.15), to that used in the recent works [26, 11]. There, instead of considering (1.1) directly, one first aims to provide a power series expansion of what is essentially the logarithmic (with $z={1\over 2}\sqrt{s}$),
$u(s):=-s{d\over ds}\log\bigg{(}e^{-s/4}\Big{\langle}e^{{1\over 2}\sqrt{s}{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}\bigg{)}.$ (4.6)
The significance of (4.6) is the characterisation of $u(s)$ as the solution of the $\sigma$-Painlevé III equation [45, 29]
$(su^{\prime\prime})^{2}+u^{\prime}(4u^{\prime}-1)(u-su^{\prime})=0$ (4.7)
with boundary condition for small $s$
$u(s)=-{s^{1+l}\over 2^{2(1+l)}\Gamma(1+l)\Gamma(2+l)}+{\rm O}(s^{2+l}).$ (4.8)
Writing $u(s)$ as a power series consistent with (4.8) and substituting in (4.7) provides a recurrence for successive coefficients. In this way, after integrating and exponentiating as required by (4.6), we have available terms in the power series expansion of the average in (4.6), and thus from (1.3) (after doing this for $l=1,2,\dots,N^{*}$) knowledge of $\{T_{l}(N)\}_{l=1,\dots,N}$ for values of $N$ up to $N^{*}$. By carrying out this procedure, tabulation up to $N^{*}=700$ was obtained in [26]. A computational improvement of this strategy was provided in [11], based on the observation that differentiating (4.7) to convert it to a particular third order nonlinear equation in the Chazy-I class. The latter is linear in the higher derivative ((4.7) is quadratic), and is quadratic in lower order terms ((4.7) is cubic). Based on this third order nonlinear equation, a tabulation up to $N^{*}=1000$ was attained in [11]. On the other hand, the recurrence (3.15) of Proposition 3.2 provides an alternative to the use of a nonlinear equation for purposes of computing the tabulation. In this formalism $T_{l}(N^{*})$ is read off from the vector $\mathbf{d}_{N^{*}}$. With the vector $\mathbf{d}_{0}$ known from (3.14), successive vectors are computed by solving the linear system in the recurrence (3.15) by back substitution. This only takes of order $l$ steps since the matrix on the right hand side is bidiagonal. With $N^{*}$ iterations the total number of steps in then of order $N^{*}l$, which for small $l$ at least thus provides a computational saving relative to the method of [11]. One remarks that the use of such tabulations was to probe correction terms to the leading scaled functional form of the distribution of the length of the longest increasing subsequence [26, 11, 12].
@@
The result of Proposition 3.2 with $\beta=2$, $n=l$, $\tilde{B}_{p}=p(2l-p)$ gives, according to (2.9) and (3.13), the matrix integral in (1.7) as proportional to $\hat{I}_{n}(x)$. Following the method of Proposition 4.2, we can similarly specify the polynomials in the scalar linear differential equation reduction of the matrix differential equation from Proposition 3.2 for small values of $l$ (without this latter restriction, the coefficients in the corresponding polynomials soon become unwieldy due to the number of digits involved and their increasing degree; recall the listing of Proposition 4.2).
@@
As revised in Appendix A, the work [36] has given an alternative way to derive the degree $l+1$ scalar linear differential equation satisfied by the matrix integral (1.7). In the case $l=2$ this is made explicit in [36, Eq. (97)], and consistency with the third order equation listed above can be checked. Implementing the recursive strategy leading to (A.12) of the Appendix in the cases $l=3,4,5$ using computer algebra establishes consistency of the results obtained by the alternative method for all the differential equations listed in Proposition 4.3.
@@
cf. (4.6). Using results from [29] it was noted that $v(s)$ satisfies the particular $\sigma$-Painlevé III equation
@@
(cf. (4.7)) subject to the small-$s$ boundary condition
@@
It was shown in [30] that this allowed for the generation of the first $l$ even powers recursively, as was the need of the application therein. However, missed in [30] was the fact made explicit in [36] that the boundary condition (4.11) does not uniquely determine a solution of (4.11), with their in fact being a family of solutions which require the specification of the coefficient of $s^{2l+1}$ to be made unique beyond degree $s^{2l}$, and which was incorrectly set to zero.
The finding in [36] of a scalar differetial equation characterisation of the matrix integral (1.7) overcame this latter point, and was used to obtain in particular the coefficients up to degree $s^{4l}$ as required for the application to the moments of the second derivative of the Riemann zeta function on the critical line as outlined in the text of the Introduction including (1.8). The present work complements such a computational objective, as in keeping with the text below (4.7) use can be made of the matrix differential equation formalism to provide efficient tabulations based on solving a vector recurrence via back substitution.
@@ Acknowledgement
of weight $m$, where $m\geq j\geq 1$. By applying [36, Thm. 3.1] to general Hankel determinants shifted by $Y_{m,j}$ and specializing to $\tau_{l,Y_{m,j},\alpha}$, we obtain the following inverse linear system.
@@
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\\ \end{pmatrix}=B^{(m)}\begin{pmatrix}\vdots\\ \sum_{h=1}^{j-1}(-1)^{h}T_{h}\tau_{l,Y_{m-h,j-h},\alpha}\\ \vdots\\ T_{m}\tau_{l,\alpha}\\ \end{pmatrix}_{j=2,\ldots,m},$
where $B^{(m)}=(b_{j,k}^{(m)})_{i,j=1,\ldots,m}$ is the $m\times m$-matrix satisfying
@@
we can express $\tau_{l,Y_{1,1},\alpha}$, $\tau_{l,Y_{2,1},\alpha}$, and $\tau_{l,Y_{2,2},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. The following lemmas for general $\alpha$ are obtained by a similar argument to that used in [36, Prop. 4.1 and Prop. 4.2] for $\alpha=1$.
@@
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\end{pmatrix}$= $-\,\sqrt{x}B^{(m)}\begin{pmatrix}\frac{d}{dx}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \frac{d}{dx}\tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}+\frac{l^{2}+l(\alpha-1)+m-1}{2\sqrt{x}}B^{(m)}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}$
$-\,\frac{1}{\sqrt{x}}C_{1}^{(m)}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\end{pmatrix}+C_{2}^{(m)}\begin{pmatrix}\tau_{l,Y_{m-2,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-2,m-2},\alpha}\end{pmatrix}$, (A.8)
where $B_{1}^{(m)}$ is given as in (A.3), $C_{1}^{(m)}=(c_{j,k}^{(1)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-1\end{subarray}}$ be an $m\times(m-1)$-matrix satisfying
@@
and $C_{2}^{(m)}=(c_{j,k}^{(2)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-2\end{subarray}}$ is an $m\times(m-2)$-matrix satisfying
@@
The above proposition generalizes [36, Prop. 4.3] from the case $\alpha=1$ to general $\alpha$. It can be rewritten as a first order matrix linear differential equation for the vector on the LHS analogous to (but distinct from) that implied by Proposition 3.2, with $n=m$, $\beta=2$ and $\tilde{B}_{p}$ given by (3.16) with $q=m+\alpha$. On this latter point, note that the entries in the corresponding vectors agree in the final entry but involve different functions otherwise.
@@
Not only $\tau_{l,Y_{l+1,l+1},\alpha}$, but also $\tau_{l,Y_{m,j},\alpha}$ for any $m,j$ with $l+1\leq j\leq m$, which can be obtained recursively from formula (A.5) as explained in [36, Remark 3.5], should be 0. Hence, this will also lead to a linear differential equation of a form similar to (A.12), but of degree $m$.

### v2

@@ 4.1 The case of general parameters
Next we would like to relate (3.10) to a scalar linear differential equation of degree $n+1$ for $I_{n}(x)$ (and thus by the replacement $b\mapsto b-1$ for $I_{0}(x)$; recall (3.12)).
@@
Let $I_{n}(x)$ be specified via the differential-difference system (3.8). For some polynomials $\{q_{m}(x)\}_{m=0,1,\dots,n-1}$ we have
$\Big(x^{n}{d^{n+1}\over dx^{n+1}}+q_{n}(x)x^{n-1}{d^{n}\over dx^{n}}+q_{n-1}(x)x^{n-2}{d^{n-1}\over dx^{n-1}}+\cdots+q_{2}(x)x{d^{2}\over dx^{2}}+q_{1}(x){d\over dx}+q_{0}(x)\Big)I_{n}(x)=0.$ (4.1)
The degree in $x$ of the coefficient of each $d^{j}\over dx^{j}$} is [(n1+j)/2][(n-1+j)/2], thus restricting the degrees of the $q_{m}(x)$. Furthermore, when $\{B_{j}\}$ in (3.8) are all integers, the coefficients of these polynomials are also integers.
@@
Let [^2] This notation is not to be confused with $D_{p}$ in (3.4). $D_{x}:=x{d\over dx}$. From (3.8) in the case $p=n$ we read off that
$nI_{n-1}(x)=(B_{n}+D_{x})I_{n}(x).$ (4.2)
@@
$n(n-1)(n-2)I_{n-3}(x)=\bigg((B_{n-2}+D_{x})\Big((B_{n-1}+D_{x})(B_{n}+D_{x})-nx\Big)-2x(n-1)(B_{n}+D_{x})\bigg)I_{n}(x).$ (4.4)
@@
$x{d\over dx}+c_{0},\>\>x^{2}{d^{2}\over dx^{2}}+d_{0}x{d\over dx}+(c_{1}+c_{2}x),\>\>x^{3}{d^{3}\over dx^{3}}+e_{0}x^{2}{d^{2}\over dx^{2}}+(d_{1}+d_{2}x)x{d\over dx}+(c_{3}+c_{4}x),$ (4.5)
for certain coefficients.
Repeating this procedure down to the case $p=1$ of (3.8) allows each $I_{j}(x)$ ($j=0,\dots,n)$ to be written as a differential operator of degree njn-j acting on $I_{n}(x)$. In particular, $n!I_{1}(x)$ has the structure of a differential operator of the form (4.1) with $n\mapsto n-1$ and an extra factor of $x$ multiplying each of the terms involving a derivative. Similarly, $n!I_{0}(x)$ has the structure of a differential operator of the form in (4.1) as written, except for an extra factor of $x$ multiplying each of the terms involving a derivative.
@@
${d\over dx}I_{0}(x)-nI_{1}(x)=0.$ (4.6)
From the established formulas for $I_{0}(x),I_{1}(x)$, this implies (4.1). That the coefficients of the polynomials $\{q_{j}(x)\}$ are integers for $\{B_{j}\}$ integers is immediate from the structure of the RHS of (4.2)–(4.4) as extends to $I_{0}(x),I_{1}(x)$. ∎
@@
In the case $\beta=2$, $n=l$ and $B_{p}=p(l-p)$, $I_{n}(x)$ is according to (2.9) and (3.13) (the latter modified in keeping with Remark 3.1.1) equal to the matrix average in (1.3) with $z=x^{1/2}$ and thus is a generating function for the sequence $\{T_{l}(N)\}_{N=1,2,\dots}$. Using different methods, the precise form of the corresponding differential operator in (4.1) has been computed for $l=2,3,\dots,7$ in [10]. From the substitution method given in the proof of Proposition 4.1 we can reclaim the listing of [10], and extend it to higher values of $l$ (see the work [45] for other applications in random matrix theory of this method in reducing first order linear matrix differential equations to higher order scalar differential equations). Here we make note of the case $l=8$.
@@
$q_{8}(x)=120,\>\>\>q_{7}(x)=12(483-10x),\>\>\>q_{6}(x)=4(36382-2541x),$
$q_{5}(x)=12(170417-27414x+364x^{2}),\>\>\>q_{4}(x)=24(214803+9830x),$
$q_{3}(x)=4(16531678-10189630x+1112127x^{2}-13120x^{3}),$
$q_{2}(x)=-8(-14762834+19053398x-4421047x^{2}+185712x^{3}),$
$q_{1}(x)=8(5883904-26926706x+13943655x^{2}-1483008x^{3}+18432x^{4}),$
$q_{0}(x)=256(-183872+380121x-98628x^{2}+4608x^{3}).$
In (1.3) let us replace $z$ by $x^{1/2}$ so that it reads
$\Big\langle e^{x^{1/2}{\rm Tr}(U+U^{\dagger})}\Big\rangle_{U(l)}=1+\sum_{N=1}^{\infty}{T_{l}(N)\over(N!)^{2}}x^{N}.$ (4.7)
As has just been illustrated in Proposition 4.2, we have from Proposition 4.1 that (4.7) satisfies a certain scalar linear differential equation of degree $l+1$. As remarked in this context in [10], it is well known and classical that one can translate the latter into a recurrence for $\{T_{l}(N)\}_{N=1,2,\dots}$ with coefficients that are polynomial in $N$. Previous literature contains the explicit form of the difference equations for small $l$; [12, footnote 22] gives references containing the explicit forms up to $l=5$. As an explicit example, from [52], for $l=4$ we have
$(N+4)(N+3)^{2}T_{4}(N)=(20N^{3}+62N^{2}+22N-24)T_{4}(N-1)-64N(N-1)^{2}T_{4}(N-2),$ (4.8)
valid for $N=1,2,\dots$. Iterating this, starting with $N=1,T_{4}(0)=1$ indeed generates the sequence (1.2). Below the proof of Proposition 4.4 note is made of the explicit forms for $l=2,3$ and 5.
From [11], one has the conjecture that $\{T_{l}(N)\}_{N=1,2,\dots}$ satisfy a linear difference equation of the form
$\sum_{k=0}^{[l/2]+1}P_{k}(N)T_{l}(N-k)=0,$ (4.9)
with each $P_{k}(N)$ a polynomial in $N$. However, already with $l=4$ we see from (4.8) that (4.9) is inaccurate, as (4.8) involves only $\{T_{4}(N),T_{4}(N-1),T_{2}(N-2)\}$ whereas (4.9) gives too a term involving $T_{2}(N-3)$. Making use of Proposition 4.1 we can (by inspection) give the corrected form of (4.9).
Proposition 4.3.
Let $\{T_{l}(N)\}_{N=1,2,\dots}$ be as in (4.7). We have that this sequence satisfies a linear difference equation of the form (4.10) with the upper terminal replaced by [(l1)/2]+1[(l-1)/2]+1.
Alternative to first determining the scalar differential equation and then converting it to a difference equation, it is possible to deduce the latter directly from the differential-difference system (3.8). [^3] We thank F. Bornemann for this observation.
Proposition 4.4.
Consider the differential-difference system (3.8). By substituting $I_{p}(x)=\sum_{N=0}^{\infty}d_{p,N}x^{N}$ as is consistent with (3.11), a difference equation for $\{d_{p,N}\}_{N=0,1,\dots}$ of the form
$\sum_{k=0}^{[(n-1)/2]+1}Q_{k}(N)d_{n,N+1-k}=0,$ (4.10)
with each $Q_{k}(p)$ a polynomial in $p$, can be arrived at.
Proof.
Substituting the power series for $I_{n-1}(x),I_{n}(x)$ in (4.2) and equating coefficients of $x^{N}$ gives
$d_{n-1,N}={1\over n}(B_{n}+N)d_{n,N}.$ (4.11)
Carrying out the analogous procedure in (3.8) with p=njp=n-j, $j=1,\dots,n-1$ gives
$d_{n-j-1,N}={1\over n-j}\Big((B_{n-j}+N)d_{n-j,N}-jd_{n-j+1,N-1}\Big).$ (4.12)
After successive substitutions, at the completion of step $j$, one has an expression for $d_{n-j-1,N}$ in terms of $\{d_{n,N^{\prime}}\}_{N^{\prime}=N-1-[j/2],\dots,N}$. In relation to $d_{0,N}$ and $d_{1,N}$ there is the additional equation
$(N+1)d_{0,N+1}-nd_{1,N}=0$ (4.13)
as follows from (4.6). Substituting the two previously obtained expressions for $d_{0,N},d_{1,N}$ in terms of $\{d_{n,N^{\prime}}\}$ we deduce (4.10). ∎
As noted in the first line of this subsection, in the present context, $n=l,B_{p}=p(l-p)$. Further, with $\{T_{l}(N)\}$ as in (4.10) we have $d_{l,N}=T_{l}(N)/(N!)^{2}$. In terms of $\{T_{l}(N)\}$, in low order cases the procedure of the proof of Proposition 4.4 gives
$(N+2)T_{2}(N+1)=(4N+2)T_{2}(N),\quad(N+3)^{2}T_{3}(N+1)=(9+22N+10N^{2})T_{3}(N)-9N^{2}T_{3}(N-1).$ (4.14)
Note that the solution of the first recurrence with $T_{2}(0)=1$ is the Catalan number $T_{2}(N)={1\over N+1}\binom{2N}{N}$; see [52] for references. The procedure reclaims (4.8) for $l=4$, and for $l=5$ gives the difference equation (4.10) with upper terminal on the sum $k=3$ and [^4] This corrects the corresponding recurrence in [52, two displayed equations below (18)], which is reported with $P_{3}(N)$ of opposite sign.
$P_{0}(N)=(N+4)^{2}(N+6)^{2},\>\>P_{1}(N)=-375+400N+843N^{2}+322N^{3}+35N^{4}$
$P_{2}(N)=(45+622N+259N^{2})(N-1)^{2},\>\>P_{3}(N)=-255(N-2)^{2}(N-1)^{2}.$ (4.15)
From the explicit forms (4.14), (4.8) and (4.2) one sees an obvious pattern in that the degree of the polynomial coefficients within each case is the same, and equal to l1l-1. However the cancellations which take place in going from the difference equation in Proposition 4.4 to the one for $\{T_{l}(N)\}$ makes this difficult to anticipate.
Remark 4.1.
1. We know that the matrix differential equation of Proposition 3.2 has a unique solution with $d_{p,0}=\delta_{p,l}\,(p=0,\dots,l)$ in (3.11); recall Remark 3.1 point 1. As a consequence it must be that the scalar differential equation (4.1) has a unique power series solution subject only the the normalisation $I_{n}(0)=1$ (i.e. that the constant term in the power series is fixed). For this latter property to hold, it must be that the corresponding difference equation (4.10) (corrected according to Proposition 4.3) must have the property that the coefficients of $T_{l}(-1),T_{l}(-2),\dots,T_{l}(-[(l-1)/2])$, whenever they appear, must vanish. Note that this property is illustrated in (4.8) and (4.2).
@@
2. In relation to computing the power series solution for general $l$, there is no advantage in first reducing the matrix differential equation down to a scalar differential equation, rather use of the analogue of (3.18) (which is based on the matrix differential equation) is more efficient. As already mentioned in Remark 3.1, point 2, a quantification of the complexity when using this approach is given in Appendix B.
@@
The result of Proposition 3.2 with $\beta=2$, $n=l$, $B_{p}=p(2l-p)$ gives, according to (2.9) and (3.13), the matrix integral in (1.7) as proportional to $I_{n}(x)$. Following the method of Proposition 4.2, we can similarly specify the polynomials in the scalar linear differential equation reduction of the matrix differential equation from Proposition 3.2 for small values of $l$ (without this latter restriction, the coefficients in the corresponding polynomials soon become unwieldy due to the number of digits involved and their increasing degree; recall the listing of Proposition 4.2).
@@
As revised in Appendix A, the work [39] has given an alternative way to derive the degree $l+1$ scalar linear differential equation satisfied by the matrix integral (1.7). In the case $l=2$ this is made explicit in [39, Eq. (97)], and consistency with the third order equation listed above can be checked. Implementing the recursive strategy leading to (A.12) of the Appendix in the cases $l=3,4,5$ using computer algebra establishes consistency of the results obtained by the alternative method for all the differential equations listed in Proposition 4.5.
@@
cf. (B.2). Using results from [32] it was noted that $v(s)$ satisfies the particular $\sigma$-Painlevé III equation
@@
(cf. (B.3)) subject to the small-$s$ boundary condition
@@
It was shown in [33] that this allowed for the generation of the first $l$ even powers recursively, as was the need of the application therein. However, missed in [33] was the fact made explicit in [39] that the boundary condition (4.18) does not uniquely determine a solution of (4.18), with there in fact being a family of solutions which require the specification of the coefficient of $s^{2l+1}$ to be made unique beyond degree $s^{2l}$, and which was incorrectly set to zero (albeit without consequence to the application given their related to the moments of the first deviative of the Riemann zeta function).
The finding in [39] of a scalar differential equation characterisation of the matrix integral (1.7) overcame this latter point, and was used to obtain in particular the coefficients up to degree $s^{4l}$ as required for the application to the moments of the second derivative of the Riemann zeta function on the critical line as outlined in the text of the Introduction including (1.8). The present work complements such a computational objective, with the vector recurrence (3.15) as deduced from the matrix differential equation and analysed from a complexity viewpoint in the second subsection of Appendix B, providing an efficient scheme for this purpose.
@@ Acknowledgement
of weight $m$, where $m\geq j\geq 1$. By applying [39, Thm. 3.1] to general Hankel determinants shifted by $Y_{m,j}$ and specializing to $\tau_{l,Y_{m,j},\alpha}$, we obtain the following inverse linear system.
@@
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\\ \end{pmatrix}=\mathbf{B}_{m}\begin{pmatrix}\vdots\\ \sum_{h=1}^{j-1}(-1)^{h}T_{h}\tau_{l,Y_{m-h,j-h},\alpha}\\ \vdots\\ T_{m}\tau_{l,\alpha}\\ \end{pmatrix}_{j=2,\ldots,m},$
where $\mathbf{B}_{m}=(b_{j,k})_{i,j=1,\ldots,m}$ is the $m\times m$-matrix satisfying
@@
we can express $\tau_{l,Y_{1,1},\alpha}$, $\tau_{l,Y_{2,1},\alpha}$, and $\tau_{l,Y_{2,2},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. The following lemmas for general $\alpha$ are obtained by a similar argument to that used in [39, Prop. 4.1 and Prop. 4.2] for $\alpha=1$.
@@
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\end{pmatrix}$= $-\,\sqrt{x}\mathbf{B}_{m}\begin{pmatrix}\frac{d}{dx}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \frac{d}{dx}\tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}+\frac{l^{2}+l(\alpha-1)+m-1}{2\sqrt{x}}\mathbf{B}_{m}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}$
$-\,\frac{1}{\sqrt{x}}\mathbf{C}^{(1)}_{m}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\end{pmatrix}+\mathbf{C}^{(2)}_{m}\begin{pmatrix}\tau_{l,Y_{m-2,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-2,m-2},\alpha}\end{pmatrix}$, (A.8)
where $\mathbf{B}_{m}$ is given as in (A.3), $\mathbf{C}^{(1)}_{m}=(c_{j,k}^{(1)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-1\end{subarray}}$ is an $m\times(m-1)$-matrix satisfying
@@
and $\mathbf{C}^{(2)}_{m}=(c_{j,k}^{(2)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-2\end{subarray}}$ is an $m\times(m-2)$-matrix satisfying
@@
The above proposition generalizes [39, Prop. 4.3] from the case $\alpha=1$ to general $\alpha$. It can be rewritten as a first order matrix linear differential equation for the vector on the LHS analogous to (but distinct from) that implied by Proposition 3.2, with $n=m$, $\beta=2$ and $B_{p}$ given by (3.16) with $q=m+\alpha$. On this latter point, note that the entries in the corresponding vectors agree in the final entry but involve different functions otherwise.
@@
Not only $\tau_{l,Y_{l+1,l+1},\alpha}$, but also $\tau_{l,Y_{m,j},\alpha}$ for any $m,j$ with $l+1\leq j\leq m$, which can be obtained recursively from formula (A.2) as explained in [39, Remark 3.5], should be 0. Hence, this will also lead to a linear differential equation of a form similar to (A.12), but of degree $m$.